# Efficient LLM Jailbreak via Adaptive Dense-to-sparse Constrained Optimization

### v1

@@ Abstract
Recent research indicates that large language models (LLMs) are susceptible to jailbreaking attacks that can generate harmful content. This paper introduces a novel token-level attack method, Adaptive Dense-to-Sparse Constrained Optimization (ADC), which effectively jailbreaks several open-source LLMs. Our approach relaxes the discrete jailbreak optimization into a continuous optimization and progressively increases the sparsity of the optimizing vectors. Consequently, our method effectively bridges the gap between discrete and continuous space optimization. Experimental results demonstrate that our method is more effective and efficient than existing token-level methods. On Harmbench, our method achieves state of the art attack success rate on seven out of eight LLMs. Code will be made available. Trigger Warning: This paper contains model behavior that can be offensive in nature.
@@ 1 Introduction
Recent advancements have allowed large language models (LLMs) to be employed across various sectors, such as content generation [15], programming support [13], and healthcare [7]. Nevertheless, LLMs can pose risks by possibly generating malicious content, including writing malware, guidance for making dangerous items, and leaking private information from their training data [18, 10]. As LLMs become more powerful and widely used, it becomes increasingly important to manage the risks associated with their misuse. In this context, the concept of red-teaming LLMs is introduced to test the reliability of their safety features [2, 17]. Consequently, the LLM jailbreak attack was developed to support the red-teaming process: by combining the jailbreak prompt with malicious questions (e.g., how to make explosives), it can mislead the aligned LLMs to circumvent the safety features and potentially produce responses that are harmful, discriminatory, violent, or sensitive.
Recently, a number of automatic jailbreak attacks have been introduced. Generally, these can be categorized into two types: prompt-level jailbreaks [8, 11, 3] and token-level jailbreaks [18, 6, 9]. Prompt-level jailbreaks employ semantically meaningful deception to compromise LLMs. These queries can be created manually, through prompt engineering, or generated by another LLM. The main advantage of prompt-level jailbreaks is that the attack query is typically in natural language, making it interoperable. However, this feature also makes these jailbreak methods vulnerable to alignment training, since training examples are easy to create [5, 1]. For instance, prompt-level jailbreaks generally show poor performance on the Llama2 series [15] (although Llama2 is not designed to defend these jailbreak methods). Token-level jailbreaks add an adversatial suffix to the user query, and use gradient information to optimize the adversatial tokens. Although token-level jailbreaks allow more precise control over the LLM output, they are more computationally expensive than prompt-level jailbreaks.
In this paper, our focus is on token-level jailbreaks. The current state of the art token-level jailbreak method, GCG [18], employs a simple coordinate descent to perform a discrete optimization in the token space. Given the inherent difficulty of discrete optimizations, finding a solution requires significant computational resources and may not always yield an optimal outcome [6].
@@
Motivated by this difficulty, our method relax the discrete optimization into continuous optimization, which allows us to employ stronger optimizers in the continuous space beyond coordinate descent. While relaxing optimization in a continuous space is straightforward, transitioning the optimizing vectors from continuous space to the discrete token space proves challenging. Direct projection onto the discrete space is ineffective, as it substantially alters the optimized vectors, leading to performance drops. To address this, we propose ADS, an adaptive dense-to-sparse constraint, on the dense token optimization. Initially, the optimization occurs in the full continuous space. As the optimization progresses and the loss decreases, we gradually reduce the optimization space to a highly sparse space, and ultimately to an approximate one-hot space. This progressive approach minimizes the gap between the relaxed contiguous space and the one-hot space and reduce the performance drop when transitioning the optimizing vectors to discrete tokens.
@@
Our approach surpasses existing token-level jailbreak methods in both effectiveness and efficiency. In an apple-to-apple comparison on AdvBench, our method significantly outperforms GCG across three LLMs with 2/3 computational cost. Additionally, on Harmbench, our method achieves state-of-the-art results for seven LLMs. Notably, our method also attains 35% attack success rate (ASR) in attacking the adversarially trained LLM, Zephyer R2D2 [10], a feat that previous token-level jailbreak methods could not achieve, as they typically yielded nearly 0% ASR performance. Codes will be available.
@@ 2.1 Problem Setup
The red string ( ! ! ! ! ! !) represents an adversarial string that causes the LLM chatbot to reply with a positive sentence, as shown in the example above. Following early work [18], we hypothesize that if the initial response of a large language model (LLM) chatbot starts with an affirmative phrase (e.g., “Sure, here is”), it is more likely to adhere to the user’s instructions. Therefore, the main objective of our jailbreak is to optimize the conditional probability of eliciting a specific target response, based on the user query, any associated system prompts, and the adversarial string.
@@ 2.2 Relaxed continuous optimization
The optimization space of Equation 2 is discrete, thus common optimizer does not work for it directly. A straightforward method is to consider a relaxed continuous optimization problem. We use the probability space as the continuation of the one-hot vector set $\mathcal{P}=\{w\in\mathbb{R}^{V}\big{|}w[i]\geq 0,\|w\|_{1}=1\}$, and relax Equation 2 to Equation 3:
@@
The optimizing loss would increase sharply after this projection. This occurs because, in most cases, the optimized result of $z_{1:n}$ from Equation 3 are dense vectors. The distance between a dense vector $x$ and its projection to $\mathcal{C}$ tends to be substantial, i.e., $\|\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(x)-x\|$ is large. The projection changes the optimizing $z_{1:n}$ greatly thus hurts optimization.
To reduce the negative impact caused by projection, we would hope that the optimizing vectors $z_{1:n}$ are sparse before we apply the projection (Equation 4). Then the distance between a sparse vector $x$ and its projection could be smaller.
To tackle this issue, we propose a dense-to-sparse constraint to the optimization of Equation 3. During the optimization process, we progressively enhance the sparsity of the optimizing vectors $z_{1:n}$, which helps decrease the loss when projecting from set $\mathcal{P}$ to set $\mathcal{C}$. Specifically, at each optimization step, we manually convert $z_{1:n}$ to be $S$-sparsity, where $S$ is given by the cross entropy loss in Equation 3:
$S=\frac{1}{m}\sum_{k=1}^{m}\exp(\text{CE loss of token }y_{k}).$ (5)
The motivation of the adaptive sparsity (Equation 5) is to use less sparsity constrain before we can find an optimal solution in the relaxed space $\mathcal{P}$. If the LLM is random guessing when predicting the target response $y_{1:m}$, the cross entropy loss is approximately $-\log(1/V)$ (recall $V$ is the vocabulary size). In this case, $S\approx V$ gives no constraint to $z_{1:n}$. If the loss in Equation 3 is close to zero, $S\approx 1$ indicates we would project all $z_{1:n}$ to the one-hot space $\mathcal{C}$ because there is no further optimization room in the relaxed space $\mathcal{P}$.
Note that an $S$-sparsity subspace in a $d$-dimensional space ($S<d$) is not a convex set. Projection onto the $S$-sparsity subspace may computationally complex and not yield a unique solution. We propose a simple transformation that convert an arbitrary vector to be $S$-sparsity and in the relaxed set $\mathcal{P}$. While this transformation is not a projection (i.e., mapping $x$ to the closest point in $\mathcal{P}$), we find it works well. Algorithm 1 describes the details of the transformation.
@@
For example suppose $S=1.2,n=20$, then $4$ random vectors from $z_{1:n}$ are converted to be 2-sparse and 16 random vectors from $z_{1:n}$ are converted to be 1-sparse, i.e., one-hot vectors. Such design is to progressively reduce the gap between the relaxed set $\mathcal{P}$ and the one-hot vector $\mathcal{C}$ especially when the loss in the relaxed space is low.
@@ 2.3 Optimizer design to escape local minima
Equation 3 possesses many local minima, most of which are suboptimal for jailbreak. Consequently, it is essential to create an optimizer designed to effectively escape local minima. We adopt the following three designs:
@@ 2.4 Complete Attack Algorithm
Algorithm 2 provides an overview of our method with one initialization start. We employ a momentum optimizer with a learning rate of 10 and a momentum of 0.99, optimizing over maximum 5000 steps.
@@
7:      Update $z_{1:n}$ using the momentum optimizer. $\triangleright$ Unconstrained optimization Convert $z_{1:n}$ to the target sparsity $\exp(\text{loss})$ using Algorithm 1. Convert $z_{1:n}$ to the target sparsity $\exp(\text{loss})$ using Algorithm 1.
8:      Convert $z_{1:n}$ to the target sparsity $\exp(\text{loss})$ using Algorithm 1. Evaluate jailbreaking with $\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(z_{1:n})$ using Equation 4. If yes, early stop. Evaluate jailbreaking with $\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(z_{1:n})$ using Equation 4. If yes, early stop. Evaluate jailbreaking with $\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(z_{1:n})$ using Equation 4. If yes, early stop.
@@ 3.1 Experimental Setup
We attempt to jailbreak both open- and closed-source LLMs. Specifically, we consider open-source LLMs: Llama2-chat-7B [15], Vicuna-v1.5-7B [4], Zephyr-7b-$\beta$ [16], and Zephyr 7B R2D2 [10] (an adversarial trained LLM to defend against attacks). We use their preset system prompts and chat templates. Results for transfer based attack on GPT-3.5 and 4 [12] will come soon.
@@
3.2 Comparison with the baseline on AdvBench
We employ GCG as our baseline for a detailed comparison using the AdvBench dataset on 3 LLMs: Llama2-chat-7B, Vicuna-v1.5-7B and Zephyr $\beta$ 7B. Recall that GCG and ADC require a differing number of attacking steps, we use 16 initialization start to align with the computing budgets of GCG with a batch size of 512. We re-implement GCG to make sure the comparison is fair.
Table 1 shows the results on 520 examples from the AdvBench harmful behaviors subset. The computing budge indicates the maximum computational cost if not considering early stop and the Wall-clock time is the average elapsed real time per sample on a single A100 machine. As shown in Table 2, our method performs significantly better than GCG on Llama2-chat-7B, and performs similarly to GCG on other LLMs since the performance is already close to 100%. However, our method still shows advantages in wall-clock time because our method is able to find a jailbreak string with fewer steps (higher early stop rate).
@@ 3.2 Comparison with the baseline on AdvBench
ADC 98.1 0.93$\times$ 11.9 78.1
We further explore our method’s capacity for precise control over the outputs of the LLM. Precise control of LLM outputs is important for multi-turn attack and better shows the search ability of the attack method. Table 2 shows the Exact Match performance on AdvBench Strings. The early stop rate is not reported because it equals to EM in the exact matching setting. We can see that ADC is better than GCG on Vicuna-v1.5-7B and Zephyr-$\beta$-7B but worse on Llama2-chat-7B. However, ADC+ consistently outperforms the baseline in terms of performance and efficiency.
@@
ADC 87.1 0.93$\times$ 7.9 min
ADC+ 97.1 0.67$\times$ 5.7 min
3.3 Comparison with existing methods on Harmbench
Existing studies on LLM jailbreak have employed various evaluation frameworks, such as testing on different subsets of AdvBench, generating responses of varying lengths and different ASR computation methods. Therefore, direct comparison with the ASR reported in these studies is not convincing. Hopefully, HarmBench [10] replicates these methods and provides a fair comparison on the HarmBench dataset. We follow their setting of chat templates, number of generated tokens and ASR computation method. Table 1 shows the comparison between our method, ADC+ and the top performance methods reported by HarmBench [10]: GCG [18], AutoPrompt (AP) [14], PAIR [3], TAP [11] and AutoDan [8].
Table 3 illustrates that although no single jailbreak method consistently outperforms others across all large language models (LLMs), our approach achieves state-of-the-art results on seven out of eight LLMs, excluding Zephyr-R2D2. Notably, our method registers more than 50% and 20% improvements in attack success rate (ASR) on LLama2-7B-chat and Qwen-7B-chat, respectively, compared to existing methods. It’s important to note that Zephyr-R2D2 has been adversarially trained to resist token-level jailbreaks, making prompt-level methods more effective in comparison. Nevertheless, our method still demonstrates significant effectiveness on Zephyr-R2D2, achieving notable ASR where other token-level jailbreak methods like GCG achieve only single-digit performance.
@@ 3.3 Comparison with existing methods on Harmbench
LLama2-7B-chat 34.5 17.0 7.5 5.5 0.5 89.5
LLama2-13B-chat 28.0 14.5 15.0 10.5 0.0 44.5
@@
Zephyr-$\beta$-7B 90.5 79.5 70.0 83.0 97.3 99.5
Zephyr-R2D2 0.0 0.0 57.5 76.5 10.5 35.0
4 Conclusion

### v2

@@ Abstract
Recent research indicates that large language models (LLMs) are susceptible to jailbreaking attacks that can generate harmful content. This paper introduces a novel token-level attack method, Adaptive Dense-to-Sparse Constrained Optimization (ADC), which has been shown to successfully jailbreak multiple open-source LLMs. Drawing inspiration from the difficulties of discrete token optimization, our method relaxes the discrete jailbreak optimization into a continuous optimization process while gradually increasing the sparsity of the optimizing vectors. This technique effectively bridges the gap between discrete and continuous space optimization. Experimental results demonstrate that our method is more effective and efficient than state-of-the-art token-level methods. On Harmbench, our approach achieves the highest attack success rate on seven out of eight LLMs compared to the latest jailbreak methods. Trigger Warning: This paper contains model behavior that can be offensive in nature.
@@ 1 Introduction
Recent advancements have enabled large language models (LLMs) to be utilized in diverse fields such as education [27], programming [6, 32], and health [23, 17]. However, these models can also pose risks by generating malicious content, including malware, instructions for creating hazardous items, and leaking sensitive information from their training data [22, 46, 25]. As LLMs become increasingly powerful and widespread, managing the risks linked to their misuse is crucial. In response, the practice of red-teaming LLMs has been introduced to evaluate their safety mechanisms [4, 40]. The LLM jailbreak attack, which emerged from this process, involves blending malicious questions (e.g., how to make explosives) with a jailbreak prompt. This strategy can deceive safety-aligned LLMs, causing them to bypass safety measures and potentially produce harmful, discriminatory, or sensitive responses [40].
Recent developments have also seen the introduction of various automatic jailbreak attacks. These attacks fall into two main categories: prompt-level jailbreaks [20, 26, 5] and token-level jailbreaks [46, 15, 24]. Prompt-level jailbreaks use semantically meaningful deception to undermine LLMs, which can be crafted manually [3], through prompt engineering [29], or generated by another LLM [5]. The key advantage of prompt-level jailbreaks is that the attack queries are in natural language, which ensures they are interoperable. However, this same characteristic makes them susceptible to alignment training, as training examples to counter these methods are relatively easy to produce [8, 3]. For example, prompt-level jailbreaks tend to perform poorly against the Llama2-chat series [36] (see Harmbench [25]), which, although not specifically designed to counter these attacks, fine-tuned with $\sim$2000 adversarial prompts.
Token-level jailbreaks, another type of jailbreak attack, directly alter the tokens of the input query to produce a specific response from the LLM. These jailbreaks allow for more precise control over the LLM’s output due to their finer-grained optimizable space compared to prompt-level jailbreaks. A representative example of this is GCG [46], which demonstrates the effectiveness of token-level jailbreaks across LLMs (see Harmbench [25]). However, this method suffers from query inefficiency. GCG requires hundreds of thousands of queries and could take up to an hour to find a successful adversarial string for some hard examples, which limits its scalability to many applications like transfer attack [46] and adversarial training [25].
In this paper, we aim to enhance the efficiency of white-box token-level jailbreaks. GCG [46] utilizes a straightforward coordinate descent to conduct discrete optimization within the token space. Due to the inherent challenges of discrete optimizations [34, 15, 42], finding a solution demands considerable computational resources and may not always result in an optimal outcome [15]. We believe that employing a more power optimization technique for the discrete token setting could significantly improve efficiency.
@@
Motivated by the challenges of discrete optimization, our method transforms discrete optimization over $V$ discrete vectors (where $V$ represents the vocabulary size) into continuous optimization in $\mathbb{R}^{V}$, allowing us to utilize more powerful optimizers beyond coordinate descent. While transitioning to continuous space is straightforward, converting the optimized vectors back to discrete token space remains challenging. Direct projection onto the discrete space [42] is ineffective, as it significantly alters the optimized vectors and leads to performance drops. To address this, we propose ADC, an Adaptive Dense-to-sparse Constraint for dense token optimization. Initially, we relax the discrete tokens into dense vectors, and the optimization takes place in the full continuous space $\mathbb{R}^{V}$. As optimization progresses and the loss decreases, we gradually constrain the optimization space to a highly sparse space. This progression is adaptive to the optimization performance, ensuring that the constraint does not significantly hinder optimization. By the end of the process, the optimization space is constrained to a nearly one-hot vector space. Our approach minimizes the gap between the relaxed continuous space and the one-hot space, reducing the performance drop when transitioning the optimized vectors to discrete tokens.
@@
Our approach surpasses existing token-level jailbreak methods in both effectiveness and efficiency. When compared fairly with the state-of-the-art token-level jailbreak method GCG [46] on AdvBench, our method significantly outperforms GCG across three LLMs: Llama2-chat [36], Vicuna [7], and Zephyr [37], with only two-thirds of the computational cost and half the wall-clock run time. Additionally, on the Harmbench jailbreak benchmark [25], our method achieves state-of-the-art results for seven LLMs, surpassing other jailbreak methods by clear margins. Notably, our method is robust against adversarially trained LLMs that resist token-level jailbreak methods. ADC attains a 26.5% attack success rate (ASR) in attacking the adversarially trained LLM Zephyer R2D2 [25], whereas previous token-level jailbreak methods [46, 20, 15, 42] yielded nearly 0% ASR performance. This result demonstrates the strength of our proposed optimization method and highlights the limitations of previous defenses against token-level jailbreaks.
2 Related Work
LLM alignment
As large language models (LLMs) continue to develop, the urgency to ensure they align with human values is growing [11, 35, 41, 43]. In response, researchers have introduced various techniques to enhance the safety alignment of LLMs. Some approaches involve gathering high-quality training data that mirror human values and using them to modify the behavior of LLMs [1, 10, 21, 36, 39]. Additionally, other strategies focus on developing training methods such as Supervised Fine-Tuning (SFT) [38, 31], Reinforcement Learning from Human Feedback (RLHF) [45, 2, 28], and adversarial training [25] to achieve alignment. Despite these efforts, aligning LLMs safely does not always prevent their potential harmful behaviors [12, 14, 19, 44].
Automated jailbreaks
Several studies have employed manual jailbreak attacks [3, 16, 40, 29] to enhance the alignment of large language models (LLMs) or for pre-deployment testing. Nevertheless, manual jailbreaks are not scalable and often suffer from a lack of diversity. Two main types of automated jailbreaks have emerged. The first type, prompt-level jailbreaks [20, 26, 5], involves creating semantically meaningful deceptions to compromise LLMs. A key benefit of this approach is that the attack queries are in natural language, facilitating compatibility. Moreover, prompt-level jailbreaks are typically black-box attacks, enabling them to target robust closed-source LLMs.
The second category, token-based jailbreaks [13, 46, 15, 24, 42], involves appending an adversarial suffix to the user’s query and optimizing these adversarial tokens using gradient information. This method allows for more precise manipulation of the LLM’s outputs. Although token-based jailbreaks generally require access to the model’s gradient information, indicating a white-box approach, they can also be adapted for black-box scenarios to some extent. GCG [46] demonstrates the potential for the transferability of jailbreak strings to closed-source LLMs like GPT.
Discrete token optimization
Deploying adversarial attacks on language models presents a unique challenge due to the discrete nature of text. While image inputs can be approximated as continuous values due to their ordered, fine-grained pixel values, text does not share this characteristic. One approach optimizes directly over discrete tokens [9, 15, 33, 46], ensuring that all updated results remain valid inputs for language models. However, this method of discrete optimization tends to be slow, as it cannot fully utilize gradient information (the update has to be from one discrete token to another). Another strategy involves optimizing in a continuous embedding space before projecting the solutions back into discrete space [18, 42, 13]. Additionally, some researchers use Langevin dynamics sampling to generate discrete prompts from continuous embeddings [30]. Above approaches involve projecting optimized soft embeddings onto their closest one-hot embeddings, ensuring that optimization occurs within the discrete token space. However, if the nearest neighbors are far from the optimized embeddings (which often happens without specific constrains), the effectiveness of the optimization significantly decreases. Our method addresses this issue by applying an adaptive dense-to-sparse constraint, ensuring that the optimized soft embeddings could be properly sparse and sufficiently close to their respective one-hot embeddings.
@@ 3.1 Problem Setup
The red string ( ! ! ! ! ! !) represents an adversarial string that causes the LLM chatbot to reply with a positive sentence, as shown in the example above. Following early work [46], we hypothesize that if the initial response of a large language model (LLM) chatbot starts with an affirmative phrase (e.g., “Sure, here is”), it is more likely to adhere to the user’s instructions. Therefore, the main objective of our jailbreak is to optimize the conditional probability of eliciting a specific target response based on the user query, any associated system prompts, and the adversarial string.
@@ 3.2 Relaxed continuous optimization
The optimization space of Equation 2 is discrete, thus common optimizers do not work for it directly. A straightforward method is to consider a relaxed continuous optimization problem. We use the probability space as the continuation of the one-hot vector set $\mathcal{P}=\{w\in\mathbb{R}^{V}\big{|}w[i]\geq 0,\|w\|_{1}=1\}$, and relax Equation 2 to Equation 3:
@@
The optimizing loss would increase sharply after this projection. This occurs because, in most cases, the optimized result of $z_{1:n}$ from Equation 3 are dense vectors. The distance between a dense vector $x$ and its projection to $\mathcal{C}$ tends to be substantial, i.e., $\|\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(x)-x\|$ is large. The projection greatly changes the optimizing $z_{1:n}$, thus hurting optimization.
To reduce the negative impact caused by projection, we would hope that the optimizing vectors $z_{1:n}$ are sparse before we apply the projection (Equation 4). Then, the distance between a sparse vector $x$ and its projection could be smaller.
Adaptive Sparsity
To tackle this issue, we propose a dense-to-sparse constraint to the optimization of Equation 3. During the optimization process, we progressively enhance the sparsity of the optimizing vectors $z_{1:n}$, which helps decrease the loss when projecting from set $\mathcal{P}$ to set $\mathcal{C}$. Specifically, at each optimization step, we manually convert $z_{1:n}$ to be $S$-sparsity, where $S$ is given by the number of wrong predictions when optimizing the classification loss in Equation 3:
[…]
The motivation of the adaptive sparsity (Equation 5) is to use less sparsity constraint before we can find an optimal solution in the relaxed space $\mathcal{P}$. If the LLM is random guessing when predicting the target response $y_{1:m}$, Equation 5 provides an exponentially growing large sparsity and gives no constraint to $z_{1:n}$. If the classification problem in Equation 3 is well optimized and all $y_{1:m}$ are correctly predicted, $S\approx 1$ indicates we would project all $z_{1:n}$ to the one-hot space $\mathcal{C}$ because there is no further optimization room in the relaxed space $\mathcal{P}$.
Note that an $S$-sparsity subspace in a $d$-dimensional space ($S<d$) is not a convex set. Projection onto the $S$-sparsity subspace may be computationally complex and not yield a unique solution. We propose a simple transformation that converts an arbitrary vector to be $S$-sparsity and in the relaxed set $\mathcal{P}$. While this transformation is not a projection (i.e., mapping $x$ to the closest point in $\mathcal{P}$), we find it works well. Algorithm 1 describes the details of the transformation.
@@
For example, suppose $S=1.2,n=20$, then $4$ random vectors from $z_{1:n}$ are converted to be 2-sparse and 16 random vectors from $z_{1:n}$ are converted to be 1-sparse, i.e., one-hot vectors. Such a design is to progressively reduce the gap between the relaxed set $\mathcal{P}$ and the one-hot vector $\mathcal{C}$, especially when the loss in the relaxed space is low.
@@ 3.3 Optimizer design to escape local minima
Equation 3 possesses many local minima, most of which are suboptimal for jailbreak. Consequently, creating an optimizer designed to escape local minima is essential. We adopt the following three designs:
@@ 3.4 Complete Attack Algorithm
Algorithm 2 provides an overview of our method with one initialization start. We employ a momentum optimizer with a learning rate of 10 and a momentum of 0.99, optimizing over maximum 5000 steps. Following GCG [46], the number of optimizable adversarial tokens is 20 for all experiments.
@@
7:      Update $z_{1:n}$ using the momentum optimizer. $\triangleright$ Unconstrained optimization Convert $z_{1:n}$ to the target sparsity in Equation 5 using Algorithm 1. Convert $z_{1:n}$ to the target sparsity in Equation 5 using Algorithm 1.
8:      Convert $z_{1:n}$ to the target sparsity in Equation 5 using Algorithm 1. Evaluate jailbreaking with $\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(z_{1:n})$ using Equation 4. If yes, early stop. Evaluate jailbreaking with $\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(z_{1:n})$ using Equation 4. If yes, early stop. Evaluate jailbreaking with $\text{proj}_{\mathcal{P}\rightarrow\mathcal{C}}(z_{1:n})$ using Equation 4. If yes, early stop.
@@ 4.1 Experimental Setup
We attempt to jailbreak both open- and closed-source LLMs. Specifically, we consider open-source LLMs: Llama2-chat-7B [36], Vicuna-v1.5-7B [7], Zephyr-7b-$\beta$ [37], and Zephyr 7B R2D2 [25] (an adversarial trained LLM to defend against attacks). We use their default system prompts and chat templates.
@@
4.2 Comparison with existing methods
Comparison with the baseline on AdvBench
We use GCG as our baseline for a detailed comparison with our method on the AdvBench dataset, focusing on three LLMs: Llama2-chat-7B, Vicuna-v1.5-7B, and Zephyr $\beta$ 7B. Both two jailbreak methods follow the same early stop criteria and utilize the same jailbreak judger (the red teaming classifier in Harmbench [25]). It is important to note that GCG and our method require different numbers of optimization steps and search widths. To measure efficiency, we consider both the computing budget and wall-clock time. The computing budget represents the maximum computational cost (including both model forward and backward) without early stopping, i.e., run the optimization until the maximum step. In contrast, the wall-clock time is the average real-time elapsed per sample on a single A100 machine. Due to early stopping, these two metrics are not perfectly correlated. To ensure a fair comparison, we maintain the same data dtype and KV-cache settings for both methods.
Table 1 presents the results on 520 examples from the AdvBench harmful behaviors subset. Our method significantly outperforms GCG on Llama2-chat-7B and performs slightly better on other LLMs, where performance is already nearly 100%. Additionally, our method shows an advantage in wall-clock time, as it can find a jailbreak string in fewer steps, indicated by a higher early stop rate. If one only aims to match the jailbreak performance of GCG, the time required by our method can be further reduced.
@@ 4.2 Comparison with existing methods
ADC 99.8 0.93$\times$ 7.9 92.1
ADC+ 98.8 0.67$\times$ 6.1 97.5
We further explore our method’s capacity for precise control over the outputs of the LLM. Precise control of LLM outputs is important for multi-turn attack and better shows the optimization ability of the attack method. Table 2 shows the Exact Match performance on AdvBench Strings. The early stop rate is not reported because it equals to EM in the exact matching setting. We can see that ADC is better than GCG on Vicuna-v1.5-7B and Zephyr-$\beta$-7B but slightly worse on Llama2-chat-7B. However, ADC+ consistently outperforms the baseline in terms of performance and efficiency.
@@
ADC 92.9 0.93$\times$ 7.1 min
ADC+ 98.4 0.67$\times$ 4.8 min
Comparison with the state of the art methods on Harmbench
Existing studies on LLM jailbreak have used various evaluation setting, including testing on different subsets of AdvBench, generating responses of varying lengths, and employing different ASR computation methods. Consequently, direct comparison with the ASR reported in these studies is not reliable. Fortunately, HarmBench [25] replicates these methods, allowing for a fair comparison on the HarmBench dataset. We adhere to their settings for chat templates, the number of generated tokens, and ASR computation methods. Table 3 presents the comparison between our method, ADC+, and the top-performing methods reported by HarmBench [25]: GCG [46], AutoPrompt (AP) [33], PAIR [5], TAP [26], and AutoDan [20]. Other jailbreak methods with lower performance are not listed. ASR numbers for these methods are sourced from HarmBench (Standard Behaviours) [25].
Table 3 illustrates that although no single jailbreak method consistently outperforms others across all large language models (LLMs), our approach achieves state-of-the-art results on seven out of eight LLMs, excluding Zephyr-R2D2. Notably, our method registers more than 50% and 20% improvements in attack success rate (ASR) on Llama2-chat and Qwen-chat, respectively, compared to existing methods. It’s important to note that Zephyr-R2D2 has been adversarially trained to resist token-level jailbreaks, making prompt-level methods more effective in comparison. Nevertheless, our method still demonstrates significant effectiveness on Zephyr-R2D2, achieving notable ASR where other token-level jailbreak methods like GCG achieve only single-digit performance.
@@
Llama2-7B-chat 34.5 17.0 7.5 5.5 0.5 92
Llama2-13B-chat 28.0 14.5 15.0 10.5 0.0 56.5
@@
Zephyr-$\beta$-7B 90.5 79.5 70.0 83.0 97.3 100
Zephyr-R2D2 0.0 0.0 57.5 76.5 10.5 26.5
Transferability
While ADC is primarily employed as a white-box jailbreak method, it can also extend to black-box jailbreak scenarios. The black-box variant of our method,, similar to GCG, seeks a transferable adversarial string by jointly optimizing across multiple examples and models. Specifically, Vicuna-v1.5-7B and Vicuna-v1.5-13B serve as surrogate models, and optimization is performed on a subset of Advbench filed by PAIR [5]. In line with GCG, we report the transfer ASR on GPT-3.5 (gpt-3.5-turbo-0125) and GPT-4 (gpt-4-0314). Table 5 compares our method with GCG and PAIR, representing token-based and template-based attacks, respectively. The results of our method are obtained by using the GPT-4 judger proposed by PAIR.
Table 4: Transfer ASR on a subset of Advbench
method GPT3.5 GPT4
GCG 87 47
PAIR 60 62
Ours 90 52
Table 5: Ablation study on the sparsity design
sparsity Vicuna Llama2
constant = 1 63.1 29.0
constant = 2 98.8 62.3
constant = 3 98.5 0.0
adaptive (ours) 99.8 96.2
Ablation Study
The adaptive sparsity design is central to our method. Table 5 compares our adaptive sparsity design with baseline methods that use constant sparsity of 1, 2, and 3. The table shows the ASR performance for Vicuna-v1.5-7B and Llama2-chat-7B on the AdvBench behavior subset. No single constant sparsity level performs optimally across both LLMs, whereas adaptive sparsity consistently outperforms all constant sparsity levels.
We also demonstrate how certain hyperparameters affect the performance of our methods. Table 7 and Table 7 present the ablation study focusing on the learning rate and the momentum of the optimizer. We report the ASR of Llama2-chat-7B and Vicuna-v1.5-7B on AdvBench Behaviours. Our method remains robust across a wide variety of learning rates. However, achieving optimal performance is critically dependent on the appropriate use of momentum. This is because the gradient at a local point may not be very helpful due to sparsity constraints in the optimization. Utilizing a large momentum can leverage historical gradient information to help minimize the loss.
Table 6: Ablation study on the learning rate
learning rate Llama2 Vicuna
0.1 83.8 92.3
1 95.0 99.8
10 (default) 96.2 99.8
100 96.2 98.8
Table 7: Ablation study on the momentum
momentum Llama2 Vicuna
0 4.0 11.9
0.5 27.7 38.1
0.9 94.2 99.8
0.99 (default) 96.2 99.8
5 Conclusion and Limitations